# DeepC4: Deep Conditional Census-Constrained Clustering for Large-scale Multitask Spatial Disaggregation of Urban Morphology

### v1

@@ 5 Conclusion and Future Work
Our proposed novel DeepC4, a deep constrained clustering model, advances the existing classical spatial disaggregation techniques using deep learning and EO data to understand the distribution of various urban morphology characteristics at large scales. Using cluster-level constraints, multitask optimization, and multilabel conditionality, we specifically demonstrated a deep learning model that can uniquely cope with the challenges of weak supervision when inferring the Rwandan urban morphology, in terms of indicators often used in the practice of disaster risk assessment. Using several built environment data and partially available building-level groundtruth, our work has also significantly contributed new insights into the existing Rwandan exposure and physical vulnerability of GEM and METEOR towards a spatial auditing of existing coarse-grained derived information. For future work, we recommend extending this study with temporal variables to understand the evolution of urban morphology. We also recommend investigating other regions with higher diversity and uniformity across labels to assess the training quality and sensitivity.
@@ Appendix A Discrete conditional probability values among urban morphology indicators
Continuous Height
Labels in meters
Discrete Target Height Classes
@@
[…]

### v2

@@ 5 Conclusion and Future Work
Our proposed novel DeepC4, a deep constrained clustering model, advances the existing classical spatial disaggregation techniques using deep learning and EO data to understand the distribution of various urban morphology characteristics at large scales. Using cluster-level constraints, multitask optimization, and multilabel conditionality, we specifically demonstrated a deep learning model that can uniquely cope with the challenges of weak supervision when inferring the Rwandan urban morphology, in terms of indicators often used in the practice of disaster risk assessment. By considering the explicit importance of well-validated census and experts’ belief systems, our approach also contributes to the overall consistency of resulting deep learning-based maps, thereby improving their relative explainability and interpretability with respect to the current understanding of end users. Using several built environment data and partially available building-level groundtruth, our work has also significantly contributed new insights into the existing Rwandan exposure and physical vulnerability of GEM and METEOR towards a spatial auditing of existing coarse-grained derived information.
For future work, we recommend extending this study with temporal variables to understand the evolution of urban morphology. We also recommend investigating other regions with higher diversity and uniformity across labels to assess the training quality and sensitivity.
@@ Appendix A Discrete conditional probability values among urban morphology indicators
Continuous Height Labels in meters Discrete Target Height Classes
@@
$\left(0,6\right)$
$\left[6,9\right)$
$\left[9,12\right)$
$\left[12,21\right)$
$\left[21,24\right)$
$\left[24,+\infty\right)$